# Lifetime measurement of the muonic atom of enriched Si isotopes

### v1

@@ Abstract
A muonic atom, composed of a negative muon and an atomic nucleus, undergoes two primary decay processes: muon-electron decay and muon nuclear capture. The branching ratio between these two processes can be determined from the measured lifetime of the muonic atom. While past researches have examined the general trend of muon capture rates across different nuclei, experimental and theoretical investigations into the isotope dependence of the lifetime remain limited.
@@ I Introduction
A muonic atom consists of a negative muon and an atomic nucleus. The muon at the lowest atomic state ( state) undergoes either - decay or muon nuclear capture. The - decay, also known as a decay-in-orbit or Michel decay, is a conversion of the muon to an electron, an electron antineutrino, and a muon neutrino, expressed as
@@
where is the lifetime of the positive muon, viz. 2.1969811(22) s [2], and is a Huff factor [3], which takes into account a reduction in the - decay rate due to the atomic binding energy. A comprehensive study of the capture rates was conducted, with most results summarized in Ref. [4]. However, these measurements focused primarily on natural abundance targets, and data on the capture rates for isotopically enriched targets are still limited. A significant isotope dependence of the muon nuclear capture rate was pointed out by some measurements in the past [5, 6, 7, 8, 9, 10, 11]. In general, the capture rate decreases with an increase in the mass number of isotopes, because of the Pauli exclusion principle on the muon capture process [12]. As the neutron number increases, more neutron orbitals are filled, making it more difficult for a proton to transform into a neutron at higher levels by muon capture.
@@
The present experiment was motivated by our previous measurement of production branching ratios of residual nuclei by muon capture on enriched Si isotopes [20]. As the branching ratio of each residual nucleus was given as the number per muon nuclear capture, the capture probability () is necessary to obtain the absolute branching ratio. The value is obtained by
The isotope dependence of for 28,29,30Si has not been known to date, and the Primakoff and Goulard-Primakoff formulas do not accurately predict isotope dependence of [5, 6, 7, 8, 9, 4]. Therefore, in the present experiment, we have measured , or equivalently, the mean lifetimes for 28,29,30Si.
@@ II Experiment
The experiment was performed at the D1 area at the muon science facility (MUSE) in the Materials and Life Science Facility (MLF) in J-PARC [21]. A proton beam from the Rapid Cycling Synchrotron (RCS) with an energy at 3 GeV irradiated a rotating graphite target to produce pions. The primary beam had a double-pulse structure with a 25 Hz repetition and a power of approximately 900 kW. The negative muon beam, a decay product of negative pions, was transported through the D-beamline, where its momentum was selected to 32 . The beam was then delivered to the D1 area, where the experimental apparatus was installed.
@@
Figure 1 shows a schematic view of the experimental setup at the D1 area. The muon beam passed through a lead collimator with a hole size of 40 mmϕ and irradiated a target. The targets were located approximately 65 mm downstream of the exit of the lead collimator. A total of ten targets was measured in the present experiment, as listed in Table 1. The main goal of the present study is to measure the lifetimes of the muonic atom of three isotopically enriched Si targets of 28,29,30Si. The enriched targets had been used in our previous study, and details of the target composition were given in Ref. [20]. The targets were in metal chippings form encapsulated in an acrylic case, as schematically illustrated in the inset of Fig. 1. The effective thickness of the targets was approximately 1.0 g/cm2. The primary source of uncertainties in the lifetime measurement in the present setup was the - decay electrons from the light elements of carbon (C), nitrogen (N), and oxygen (O), as discussed in the next section. The silver plates were used to avoid stopping the muon beam in the acrylic case. A silver mask with a hole size of 12 mmϕ and a thickness of 0.5 mm was placed on the upstream side of the case, and a silver backing plate with a thickness of 0.5 mm was installed on the downstream side of the samples. To evaluate the effect of the chipping form of the target in the acrylic case, natural abundance silicon (natSi) samples of a self-supported metal plate and a powder form in the same acrylic case were also measured. In addition to the Si targets, the lifetimes of the muonic atoms of natMg and 27Al, which have a similar lifetime to silicon isotopes, were measured to evaluate the reliability of the measurement. For 27Al, both a self-supported plate and a plate in the acrylic case were measured to evaluate the effect of the target case. In a natAg sample measurement, the lifetimes of background components of the measurements were evaluated. The 55Mn target was measured to evaluate the reliability of lifetime measurement in the range between the lifetime of silicon isotopes and silver. Those targets had an energy loss equivalent thickness of the enriched Si targets and were used to verify the validity of the experimental and analysis procedures.
@@
The lifetime of the muonic atom was measured by detecting the decay electrons using a SR spectrometer based on Kalliope detectors [22]. The spectrometer consisted of a total of 40 units of the Kalliope detectors. Each unit of the Kalliope detectors was comprised of 16 pairs of plastic scintillators. A coincidence of each pair defined an electron event of the - decay. A zero-field correction was applied to avoid the muon spin relaxation and rotation affecting the measured lifetime. The residual magnetic field at the center of the spectrometer was less than 10 mG. The timings of the electron event were stored from to 57 s with respect to the timing of the beam arrival. The measurement times for each target are listed in Table 1.
@@ III Analysis
Backgrounds in the spectra were evaluated using an Ag metal plate target in the acrylic case, as shown in Fig. 2. Due to the double pulse structure of the beam, two peaks were observed in the timing spectra, along with two peaks at an earlier timing than the muon beam, where scattered electrons contaminated the beam. The time origin of the spectrum is defined as approximately 40 ns after the end timing of the second pulse. The spectrum was fitted using the following function consisting of three exponentials:
where is the lifetime of the natAg muonic atom, is the lifetime of the muonic atom of air and the acrylic case, is a long-lived background, and () are coefficients. The lifetimes of these three components were well separated and could be deduced by the free fitting. The second term originated from the events where the muon stopped in the air or the acrylic case. The air and acrylic contain carbon (C), nitrogen (N), and oxygen (O). Because their lifetimes are similar ( ns, ns, and ns [4]), the decay curve was represented with a single exponential with an averaged lifetime (). The value was determined to be 1.95(8) s. The quoted uncertainty contained the statistical uncertainty from the fitting (0.03 s) and the systematic uncertainty. The systematic uncertainty originates from variations in the ratio of muons stopped in the air and acrylic case, which depends on the shape of the target. Hence, an additional systematic uncertainty of 0.05 s was estimated in , which is used in subsequent analyses. The third term of Eq. (5) represents the long-lived background which was not related to the muon decay. Although the existence of such backgrounds has been known for a long time, its origin has remained unclear. It might be attributable to neutrons from the primary beam dump. The value was deduced from the fitting to be 25.9(4) s. The shape (decay constant) of the long-lived background was independent of the targets, so we used the same value for all targets. Since the events before the beam arrival were negligible, a constant term was not included in the fitting function. The lifetime of the muonic Ag was deduced to be 87.0(6) ns.
@@
where is the lifetime of the muonic atom of the target measured. Coefficients, (), are the parameters to be fitted with . If the target was a self-supported metal plate without using the Ag masks and backing plate, the decay component of Ag was omitted in the fitting, i.e., . For the self-supported metal targets, was deduced independently from the Ag sample measurement as discussed above. The primary background contribution, the second term of Eq. (5), was attributed to air, which contains less carbon. The lifetime of 1.88(5) s was used based on the composition of air and the muonic lifetime of nitrogen and oxygen. The muon stopping probability in the air was estimated to be approximately 1% using the GEANT4 toolkit [23, 24, 25]. This value was consistent with the stopping rate deduced from fits to data obtained with the self-supported metal targets. Note that the hyperfine splitting of muon state of 27Al can affect the lifetime of the muonic 27Al. Nevertheless, this effect was not considered in the present analysis, and the decay curve was fitted with a single exponential function.
@@ IV Results
| Target | Time (h) | Huff factor | Lifetime (ns) | Capture rate (s-1) | ||
| Nucleus | Form | Present | Suzuki et al.[4] | Present | ||
| 28Si | Powder in case | |||||
| 29Si | Powder in case | |||||
| 30Si | Small pieces in case | |||||
| natSi | Self-supported plate | |||||
| natSi | Powder in case | |||||
| natMg | Self-supported plate | |||||
| 27Al | Self-supported plate | |||||
| 27Al | Plate in case | |||||
| 55Mn | Chippings in case | |||||
| natAg | Plate in case | |||||
Table 1 summarized the measured lifetime and corresponding capture rates, calculated using Eq. (3). The Huff factors used in the conversion were taken from Ref. [4] and are included in the table. In this study, the lifetimes of isotopically enriched 28,29,30Si were obtained for the first time. The lifetimes of natMg, 27Al, 55Mn, and natAg were found to be consistent with previous measurement [4], confirming the reliability of our experimental setup and data analysis. However, the lifetime of natSi deviated from the previously reported value. To investigate the effect of the different forms of the targets and the acrylic case, we measured two natSi targets: a self-supported metal plate and a powder form encapsulated in the acrylic case. Both targets yielded consistent results, confirming that the shapes of the target did not influence the measured lifetime. Additionally, the lifetime of natSi was also deduced from that of the enriched targets obtained in the present study. A Monte-Carlo calculation was used to create the decay curve of natSi by summing the timing spectrum of enriched Si, weighted by the natural abundance [26]. By fitting this simulated decay curve, the average lifetime of natSi was deduced to be 750.4(42) ns, consistent with our measurements but deviated from the previously reported value. The reason for the discrepancy remains unclear.
@@ V Discussions
In this study, we have obtained the muon nuclear capture rates of ten targets. As the present study was the first investigation of muonic lifetime using the SR spectrometer at MLF, we first discuss the measurement precision. The statistical uncertainty for all targets was approximately 1–2 ns. For targets in acrylic cases, the propagated uncertainty from had a major contribution to the total uncertainty. In the specific case of 55Mn, the short lifetime of 230.7(53) ns led to additional uncertainty from the overlap with the decay curve originating from Ag, the lifetime of which is 87.0(6) ns. For the uncertainty of self-supported metal targets, namely natSi and 27Al, the statistical uncertainty dominated and the uncertainty propagated from and had a minor impact on the overall precision because of the pure sample material. The present results demonstrate the potential of our method to measure the lifetimes of various nuclei with a precision of 1–2 ns within a few hours of measurement.
Table 2 compares the capture rate obtained in the present study with theoretical calculations. The capture rates for natSi and 27Al in the table were calculated by taking an uncertainty-weighted average of two experimental data. The isotope dependence of the capture rate for the silicon isotopes is illustrated in Fig. 4. The data for natSi is also plotted at an average mass number of 28.11. The experimental capture rates show a smooth and monotonic decrease with increasing mass numbers. The smooth trend of the capture rate indicates that it is not influenced by the odd-even effect of the neutron number.
@@
where represents an overlap of the wave functions between the muon and nucleus, and and are the mass number and the atomic number of the target nucleus, respectively. While most previous studies and applications [4, 27, 28, 25, 14] have taken the values from original work by Ford and Wills [29], we use obtained by a theoretical calculation explained below. The parameters and are coefficients obtained from a global fit of the experimental data, where and taken from Ref. [4] were used in the following calculation. Equation 7 was later modified to introduce higher order corrections by Goulard and Primakoff [13] as
@@
The capture rate is predominantly influenced by two effects: an overlap of the wave functions between the muon and the nucleus, and the Pauli exclusion effect, which is expressed as , and parameters of and in the Primakoff and the Goulard-Primakoff formulas, respectively.
@@
Table 2 lists the calculated values. For the silicon isotopes, the calculated is slightly larger than the original value of [29]. Because this constant value was used to deduce parameters of and in the Primakoff and the Goulard-Primakoff formulas [4], respectively, absolute values of the calculated capture rates by these formulas may be changed by reevaluating the parameters using our values. Thus, the isotope dependence of the capture rate is discussed with relative values normalized to that of 28Si. While the reductions of are 9.4(11)% and 19.9(9)% for 29Si and 30Si, respectively, the corresponding reductions in are two order of magnitude smaller of 0.07% and 0.14%. Thus, the effect of on the isotope dependence of is negligible for the Si isotopes, in contrast to an Eu isotope case [8]. The significant isotope effect from between 151Eu88 and 153Eu90 might be due to the large change in nuclear deformation in the region of the neutron number from 82 to 92. In the case of the Si isotopes, the quadrupole deformation parameters () were experimentally deduced from the transition probability between and to be 0.407(7) and 0.315(7) for 28Si and 30Si, respectively [40]. The nuclear deformation affects through a root-mean-squere charge radius () of through Eq. (9). The experimenatal were 3.1224(24) and 3.1336(40) fm in Ref. [41] and our calculated of were 3.1469 and 3.1623 fm for 28Si and 30Si, respectively. The calculated is consistent with the experimental values, although the calculation assumed spherical symmetry. Therefore, we conclude that the isotope dependence of for the Si isotopes is not explained by that of .
The Pauli exclusion effect was expressed in the term with the parameter and in the Primakoff and the Goulard-Primakoff formulas, respectively. While the absolute capture rates calculated by the Primakoff formula underpredict experimental data, the isotope dependence is reasonably well reproduced, as shown in Fig. 4. The Goulard-Primakoff formula reproduces the capture rates of 28Si and natSi but fails to predict those of 29Si and 30Si. The inadequacy of both Primakoff and Goulard-Primakoff formulas in accounting for the isotope effect has also been previously pointed out in Refs. [4, 14]. A re-evaluation of these Primakoff’s formulas is a worthwhile direction for future research, and the experimental method established in this study offers an opportunity for systematic measurements.
@@
The measured isotope dependence of the muon capture rate was also compared with a microscopic and evaporation model (MEM) calculation [15]. In this model, the wave function of the muonic atom was calculated using the same method as that used to calculate . The muon capture rates were calculated using the second Tamm-Dancoff approximation [42] with effective interactions of SkO’ [43] and SGII [33]. Calculated capture rates strongly depend on the axial vector and induced pseudo-scaler coupling constants ( and , respectively), as well as the choice of effective interaction. As the capture rate calculated by MEM has large ambiguity due to the uncertainties in and in medium, possible combinations of the parameter set of and were deduced to reproduce the measured capture rates. Note that the values of and presented in this study are at zero momentum transfer, namely and , respectively. Due to the limitation of the MEM calculation to even-even nuclei, we discuss 28Si and 30Si hereinafter.

### v2

# Lifetime measurement of the muonic atoms of enriched Si isotopes

@@ Abstract
A muonic atom, composed of a negative muon and an atomic nucleus, undergoes two primary decay processes: muon decay and muon nuclear capture. The branching ratio between these two processes can be determined from the measured lifetime of the muonic atom. While past researches have examined the general trend of muon capture rates across different elements, experimental and theoretical investigations into the isotope dependence of the lifetime remain limited.
@@ I Introduction
A muonic atom consists of a negative muon and an atomic nucleus. The muon at the lowest atomic state ( state) undergoes either muon decay or muon nuclear capture. The muon decay, also known as a decay-in-orbit or Michel decay, is a conversion of the muon to an electron, an electron antineutrino, and a muon neutrino, expressed as
@@
where is the lifetime of the positive muon, viz. 2.1969811(22) s [2], and is a Huff factor [3], which takes into account a reduction in the muon decay rate due to the atomic binding energy. A comprehensive study of the capture rates was conducted, with most results summarized in Ref. [4]. However, these measurements focused primarily on natural abundance targets, and data on the capture rates for isotopically enriched targets are still limited. A significant isotope dependence of the muon nuclear capture rate was pointed out by some measurements in the past [5, 6, 7, 8, 9, 10, 11]. In general, the capture rate decreases with an increase in the mass number of isotopes, because of the Pauli exclusion principle on the muon capture process [12]. As the neutron number increases, more neutron orbitals are filled, making it more difficult for a proton to transform into a neutron at higher levels by muon capture.
@@
The present experiment was motivated by our previous measurement of production branching ratios of residual nuclei by muon capture on enriched Si isotopes [20]. As the branching ratio of each residual nucleus was given as the number per muon nuclear capture, the capture probability () is necessary to obtain the absolute branching ratio. The isotope dependence of for 28,29,30Si has not been known to date, and the Primakoff and Goulard-Primakoff formulas do not accurately predict isotope dependence of [5, 6, 7, 8, 9, 4]. Therefore, in the present experiment, we have measured , or equivalently, the mean lifetimes for 28,29,30Si.
@@ II Experiment
The experiment was performed at the D1 area at the muon science facility (MUSE) in the Materials and Life Science Facility (MLF) in J-PARC [21]. A proton beam from the Rapid Cycling Synchrotron (RCS) with an energy at 3 GeV irradiated a rotating graphite target to produce pions. The primary beam had a double-pulse structure with a 25 Hz repetition and a power of approximately 900 kW. The negative muon beam, a decay product of negative pions, was transported through the D-beamline. The beam momentum was selected to 32 to maximize the number of muons stopped at the target and minimize that at the air and target case. The beam was then delivered to the D1 area, where the experimental apparatus was installed. The muon beam intensity was approximately sec-1.
@@
Figure 1 shows a schematic view of the experimental setup at the D1 area. The muon beam passed through a lead collimator with a hole size of 40 mm in diameter and irradiated a target. The targets were located approximately 65 mm downstream of the exit of the lead collimator. A total of ten targets was measured in the present experiment, as listed in Table 1. The main goal of the present study is to measure the lifetimes of the muonic atom of three isotopically enriched Si targets of 28,29,30Si. The enriched targets had been used in our previous study, and details of the target form, size, weight, chemical purity, and target enrichment are given in Table 1. The targets were in metal chippings form encapsulated in an acrylic case, as schematically illustrated in the inset of Fig. 1. The effective thickness of the targets was approximately 1.0 g/cm2. The primary source of uncertainties in the lifetime measurement in the present setup was the muon decay electrons from the light elements of carbon (C), nitrogen (N), and oxygen (O), as discussed in the next section. The silver plates were used to avoid stopping the muon beam in the acrylic case. A silver mask with a hole size of 12 mm in diameter and a thickness of 0.5 mm was placed on the upstream side of the case, and a silver backing plate with a thickness of 0.5 mm was installed on the downstream side of the samples. To evaluate the effect of the chipping form of the target in the acrylic case, natural abundance silicon (natSi) samples of a self-supported metal plate and a powder form in the same acrylic case were also measured. In addition to the Si targets, the lifetimes of the muonic atoms of natMg and 27Al, which have a similar lifetime to silicon isotopes, were measured to evaluate the reliability of the measurement. For 27Al, both a self-supported plate and a plate in the acrylic case were measured to evaluate the effect of the target case. In a natAg sample measurement, the lifetimes of background components of the measurements were evaluated. The 55Mn target was measured to evaluate the reliability of lifetime measurement in the range between the lifetime of silicon isotopes and silver. Those targets had an energy loss equivalent thickness of the enriched Si targets and were used to verify the validity of the experimental and analysis procedures.
@@
The lifetime of the muonic atom was measured by detecting the decay electrons using a SR spectrometer based on Kalliope detectors [22]. The spectrometer consisted of a total of 40 units of the Kalliope detectors. Each unit of the Kalliope detectors was comprised of 16 pairs of plastic scintillators. A coincidence of each pair defined an electron event of the muon decay. A zero-field correction was applied to avoid the muon spin relaxation and rotation affecting the measured lifetime. The residual magnetic field at the center of the spectrometer was less than 10 mG. The data acquisition system was triggered by the reference timing of the synchrotron. As the Kalliope detector system can obtain signals until 1 Mcps without saturation [23], the dead time was negligible in the present experiment with a negative muon beam. The timings of the electron event were stored from to 57 s with respect to the timing of the beam arrival [24]. The measurement times for each target are listed in Table 1.
@@ III Analysis
Backgrounds in the spectra were evaluated using an Ag metal plate target in the acrylic case, as shown in Fig. 2. Due to the double pulse structure of the beam, two peaks were observed in the timing spectra, along with two peaks at an earlier timing than the muon beam, where electrons contaminated the beam scattered at the target and surrounding materials. The time origin of the spectrum is defined as approximately 40 ns after the end timing of the second pulse. The spectrum was fitted by log-likelihood method using the following function consisting of three exponentials:
where is the lifetime of the natAg muonic atom, is the lifetime of the muonic atom of air and the acrylic case, is a long-lived background, and () are coefficients. The lifetimes of these three components were well separated and could be deduced by the free fitting. The second term originated from the events where the muon stopped in the air or the acrylic case, which constitutes approximately 2% of the total count. The air and acrylic contain carbon (C), nitrogen (N), and oxygen (O). Because their lifetimes are similar ( ns, ns, and ns [4]), the decay curve was represented with a single exponential with an averaged lifetime (). The value was determined to be 1.94(8) s. The quoted uncertainty contained the statistical uncertainty from the fitting (0.05 s) and the systematic uncertainty. The systematic uncertainty originates from variations in the ratio of muons stopped in the air and acrylic case, which depends on the shape of the target. Hence, an additional systematic uncertainty of 0.03 s was estimated in , which is used in subsequent analyses. The third term of Eq. (4) represents the long-lived background which was not related to the muon decay. Although the existence of such backgrounds has been known for a long time, its origin has remained unclear. It might be attributable to neutrons from the primary beam dump. The value was deduced from the fitting to be 25.8(16) s. The shape (decay constant) of the long-lived background was independent of the targets, so we used the same value for all targets. Since the events before the beam arrival were negligible, a constant term was not included in the fitting function. The lifetime of the muonic Ag was deduced to be 87.0(5) ns.
@@
where is the lifetime of the muonic atom of the target measured. Coefficients, (), are the parameters to be fitted with . If the target was a self-supported metal plate without using the Ag masks and backing plate, the decay component of Ag was omitted in the fitting, i.e., . For the self-supported metal targets, was deduced independently from the Ag sample measurement as discussed above. The primary background contribution, the second term of Eq. (4), was attributed to air, which contains less carbon. The lifetime of 1.88(5) s was used based on the composition of air and the muonic lifetime of nitrogen and oxygen. The muon stopping probability in the air was estimated to be approximately 1% using the GEANT4 toolkit [25, 26, 27]. This value was consistent with the stopping rate deduced from fits to data obtained with the self-supported metal targets. Note that the hyperfine splitting of muon state of 27Al can affect the lifetime of the muonic 27Al. Nevertheless, this effect was not considered; the decay curve was fitted with a single exponential function following the same analysis procedure as in Ref. [4].
@@ IV Results
| Target | Time (h) | Huff factor | Lifetime (ns) | Capture rate (s-1) | ||||
| Nucleus | Form | Size (mm) | Weight (g) | Present | Suzuki et al.[4] | Present | ||
| 28Si | Powder in case | 15.02.8 | 0.500 | |||||
| 29Si | Powder in case | 15.02.8 | 0.500 | |||||
| 30Si | Small pieces in case | 15.02.8 | 0.500 | |||||
| natSi | Self-supported plate | 501.96 | 9.00 | |||||
| natSi | Powder in case | 15.02.8 | 0.550 | |||||
| natMg | Self-supported plate | 50502.0 | 8.54 | |||||
| 27Al | Self-supported plate | 50502.0 | 13.6 | |||||
| 27Al | Plate in case | 14.81.0 | 0.451 | |||||
| 55Mn | Chippings in case | 15.02.0 | 2.127 | |||||
| natAg | Plate in case | 14.80.3 | 0.544 | |||||
Table 1 summarized the measured lifetime and corresponding capture rates, calculated using Eq. (3). The Huff factors used in the conversion were taken from Ref. [4] and are included in the table. In this study, the lifetimes of isotopically enriched 28,29,30Si were obtained for the first time. The lifetimes of natMg, 27Al, 55Mn, and natAg were found to be consistent with previous measurement [4], confirming the reliability of our experimental setup and data analysis. However, the lifetime of natSi deviated from the previously reported value. To investigate the effect of the different forms of the targets and the acrylic case, we measured two natSi targets: a self-supported metal plate and a powder form encapsulated in the acrylic case. Both targets yielded consistent results, confirming that the shapes of the targets did not influence the measured lifetime. Additionally, the lifetime of natSi was also deduced from that of the enriched targets obtained in the present study. A Monte-Carlo calculation was used to create the decay curve of natSi by summing the timing spectrum of enriched Si, weighted by the natural abundance [28]. By fitting this simulated decay curve, the average lifetime of natSi was deduced to be 749.7(48) ns, consistent with our measurements as illustrated in Fig. 4. While our adopted value of 751.4(10) ns, an uncertainty-weighted average of the present results, deviates from the previous value of 756.0(10) ns, the reason for this discrepancy remains unclear.
@@ V Discussions
| Target | Lifetime | Propagated uncertainty | Total | ||||
| Range | |||||||
| 28Si | 743.9 | 2.01 | 0.44 | 4.61 | 0.85 | 0.32 | 5.0 |
| 29Si | 794.2 | 1.85 | 0.39 | 2.89 | 0.91 | 0.35 | 3.5 |
| 30Si | 859.0 | 1.96 | 0.43 | 1.62 | 1.08 | 0.16 | 2.7 |
| natSi (plate) | 751.6 | 0.77 | - | 0.68 | 0.20 | 0.20 | 1.0 |
| natSi (in case) | 749.8 | 2.26 | 0.35 | 1.40 | 0.80 | 0.18 | 2.8 |
| natMg | 1065.5 | 1.96 | - | 0.39 | 0.27 | 0.36 | 2.0 |
| 27Al (plate) | 863.7 | 0.94 | - | 0.04 | 0.23 | 0.10 | 1.0 |
| 27Al (in case) | 862.5 | 3.19 | 0.43 | 7.58 | 0.98 | 0.34 | 8.1 |
| 55Mn | 230.3 | 3.57 | 0.66 | 2.54 | 0.42 | 0.96 | 4.5 |
In this study, we have obtained the muon nuclear capture rates of ten targets. As the present study was the first investigation of muonic lifetime using the SR spectrometer at MLF, we first discuss the measurement precision. Table 2 details the systematic breakdown of the uncertainty. Propagated uncertainties from each parameter were estimated using the Monte-Carlo method by keeping other parameters fixed. The statistical uncertainty on for most targets was approximately 1–2 ns. For targets in acrylic cases, the propagated uncertainty from dominated the total uncertainty. For the self-supported metal plate targets, namely natSi, natMg, and 27Al, the statistical uncertainty dominated the total uncertainty, with uncertainties propagated from and having a minor impact on the overall precision. The present results demonstrate the potential of our method to measure the lifetimes of various nuclei with a precision of 1–2 ns within a few hours of measurement.
Table 3 compares the capture rate obtained in the present study with theoretical calculations. The capture rates for natSi and 27Al in the table were calculated by taking an uncertainty-weighted average of two experimental data. The isotope dependence of the capture rate for the silicon isotopes is illustrated in Fig. 5. The data for natSi is also plotted at an average mass number of 28.11. The experimental capture rates decrease linearly with increasing mass numbers. This constant rate of decrease indicates that the capture rate is not influenced by the odd-even effect of the neutron number.
@@
where represents an overlap of the wave functions between the muon and nucleus, and and are the mass number and the atomic number of the target nucleus, respectively. While most previous studies and applications [4, 29, 30, 27, 14] have taken the values from original work by Ford and Wills [31], we use obtained by a theoretical calculation explained below. The parameters and are coefficients obtained from a global fit of the experimental data, where and taken from Ref. [4] were used in the following calculation. Equation (6) was later modified to introduce higher order corrections by Goulard and Primakoff [13] as
@@
The capture rate is predominantly influenced by two effects: an overlap of the wave functions between the muon and the nucleus, and the Pauli exclusion effect, which is expressed as , and parameters of and in the Primakoff and Goulard-Primakoff formulas, respectively.
@@
Table 3 lists the calculated values. For the silicon isotopes, the calculated is slightly larger than the original value of [31]. Because this constant value was used to deduce parameters of and in the Primakoff and Goulard-Primakoff formulas [4], respectively, absolute values of the calculated capture rates by these formulas may be changed by reevaluating the parameters using our values. Thus, the isotope dependence of the capture rate is discussed with relative values normalized to that of 28Si. While the reductions of are 9.54(12)% and 20.18(23)% for 29Si and 30Si, respectively, the corresponding reductions in are two order of magnitude smaller of 0.10% and 0.22%. Thus, the effect of on the isotope dependence of is negligible for the Si isotopes, in contrast to an Eu isotope case [8]. The significant isotope effect from between 151Eu88 and 153Eu90 might be due to the large change in nuclear deformation in the region of the neutron number from 82 to 92. In the case of the Si isotopes, the quadrupole deformation parameters () were experimentally deduced from the transition probability between and to be 0.407(7) and 0.315(7) for 28Si and 30Si, respectively [42]. The nuclear deformation affects through a root-mean-squere charge radius () of through Eq. (8). The experimenatal were 3.1224(24) and 3.1336(40) fm in Ref. [43] and our calculated of were 3.15 and 3.16 fm for 28Si and 30Si, respectively. The calculated is consistent with the experimental values, although the calculation assumed spherical symmetry. Therefore, we conclude that the isotope dependence of for the Si isotopes is not explained by that of .
The Pauli exclusion effect was expressed in the term with the parameter and in the Primakoff and Goulard-Primakoff formulas, respectively. While the absolute capture rates calculated by the Primakoff formula underpredict experimental data, the isotope dependence is reasonably well reproduced, as shown in Fig. 5. The Goulard-Primakoff formula overestimates the capture rates of 28Si and natSi, while it underestimates those of 29Si and 30Si. The inadequacy of both Primakoff and Goulard-Primakoff formulas in accounting for the isotope effect has also been previously pointed out in Refs. [4, 14]. A re-evaluation of these Primakoff’s formulas is a worthwhile direction for future research, and the experimental method established in this study offers an opportunity for systematic measurements.
@@
The measured isotope dependence of the muon capture rate was also compared with a microscopic and evaporation model (MEM) calculation [15]. In this model, the wave function of the muonic atom was calculated using the same method as that used to calculate . The muon capture rates were calculated using the second Tamm-Dancoff approximation [44] with effective interactions of SkO’ [45] and SGII [35]. Calculated capture rates strongly depend on the axial vector and induced pseudo-scalar coupling constants ( and , respectively), as well as the choice of effective interaction. As the capture rate calculated by MEM has large ambiguity due to the uncertainties in and in medium, possible combinations of the parameter set of and were deduced to reproduce the measured capture rates. Note that the values of and presented in this study are at zero momentum transfer, namely and , respectively. Due to the limitation of the MEM calculation to even-even nuclei, we discuss 28Si and 30Si hereinafter.